OpenStudy (anonymous):

Solve the equation: 12(2y-9) = 6y(y-17)

OpenStudy (anonymous):

multiply it all out. we get 24y-108=6y^2-102y put all to one side and combine like terms. 6y^2-126y+108=0. divide by 6. y^2+21y+18=0 this becomes (y+9)(y+2)=0 so, y=-9, -2

OpenStudy (anonymous):

ok so i will do this for multiplying a polynomial by a monmial?

OpenStudy (anonymous):

Doesn't matter.

OpenStudy (anonymous):

Just distribute the terms and combine them. Then, you can find the roots by equalling it to 0.

OpenStudy (anonymous):

This will always have the form ay^2+by+c=0. a, b, and c are constants that can be negative, 0, or positive.